Q3. Prove that the square of the distance between the two points (x1 , y1) and    (x2 , y2) of the circle x2 +y2 = a2 is 2(a2-x1x2-y1y2).

Grade:12

2 Answers

APURV GOEL
12 years ago

We have to find (x1-x2)2 + (y1-y2)2 = x12 + x22 + y12 + y22 - 2x1x2 - 2y1y2 = 2a2 - 2x1x2 - 2y1y2

as x12 + y12 = a2

andx22 + y22 = a2

APURV GOEL
12 years ago

Please tell me why did u disapproved my solution.